Find the binding energy of helium 3 and helium 4
1
Expert's answer
2020-04-15T10:24:58-0400

The binding energy of 23He is


"E_b=\Delta mc^2,\\
\Delta m=Zm_p+(A-Z)m_n-m_\text{He},\\
E_b=(Zm_p+(A-Z)m_n-m_\text{He})c^2=\\
=(2\cdot 1.007 276+(3-2)\cdot1.008 664-\\-3.016029)\text{u}\cdot c^2 \cdot931.49\space\frac{\text{MeV}/c^2}{\text{u}}=6.695\text{ MeV}."

The binding energy of 24He is, likewise:


"E_b=\\=(2\cdot 1.007 276+(4-2)\cdot1.008 664-4.002603)\text{u}×\\× c^2 \cdot931.49\space\frac{\text{MeV}/c^2}{\text{u}}=27.27\text{ MeV}."